OpenStudy (anonymous):

The distance that an object will fall in (t) seconds varies directly with the square of (t). An object falls 5 feet in 1 second. How long will it take to fall 125 feet?

OpenStudy (anonymous):

If a term varies directly with t, velocity is an example, \[ v \propto t \] we can say that \[ \frac{v_!}{v_2} \propto \frac{t_!}{t_2} \] and set up the proportional equation \[ \frac{v_!}{v_2} = \frac{t_!}{t_2} \] Since x varies with t squared \[ x \propto t_2\] ....

OpenStudy (anonymous):

~~~ \[ x \propto t^2 \]

OpenStudy (anonymous):

How might you set up a proportional equation to solve for an unknown value of t?

OpenStudy (anonymous):

5/125=1/t^2 => t=5s

OpenStudy (anonymous):

\[(125 feet)\frac{ 1second }{ 5 feet}=25 seconds\] We must set up a proportion so that the object takes 1 second to fall 5 feet. Feet would cancel out leaving seconds. I could be incorrect, please let me know if I am wrong.

OpenStudy (anonymous):

@anthonykanow It's proportional to t^2, so should look like \[ \frac{x_1}{x_2} = \frac{t_1^2}{t_2^2} \] \[ 125 ft \frac {1s^2}{5ft} =25s^2 \] \[t_2^2 = 25 s^2\] So the time would be the square root of 25